# DanceGRPO: Unleashing GRPO on Visual Generation

### v2

@@ Abstract
Recent breakthroughs in generative models—particularly diffusion models and rectified flows—have revolutionized visual content creation, yet aligning model outputs with human preferences remains a critical challenge. Existing reinforcement learning (RL)-based methods for visual generation face critical limitations: incompatibility with modern Ordinary Differential Equations (ODEs)-based sampling paradigms, instability in large-scale training, and lack of validation for video generation. This paper introduces DanceGRPO, the first unified framework to adapt Group Relative Policy Optimization (GRPO) to visual generation paradigms, unleashing one unified RL algorithm across two generative paradigms (diffusion models and rectified flows), three tasks (text-to-image, text-to-video, image-to-video), four foundation models (Stable Diffusion, HunyuanVideo, FLUX, SkyReels-I2V), and five reward models (image/video aesthetics, text-image alignment, video motion quality, and binary reward). To our knowledge, DanceGRPO is the first RL-based unified framework capable of seamless adaptation across diverse generative paradigms, tasks, foundational models, and reward models. DanceGRPO demonstrates consistent and substantial improvements, which outperform baselines by up to 181% on benchmarks such as HPS-v2.1, CLIP Score, VideoAlign, and GenEval. Notably, DanceGRPO not only can stabilize policy optimization for complex video generation, but also enables generative policy to better capture denoising trajectories for Best-of-N inference scaling and learn from sparse binary feedback. Our results establish DanceGRPO as a robust and versatile solution for scaling Reinforcement Learning from Human Feedback (RLHF) tasks in visual generation, offering new insights into harmonizing reinforcement learning and visual synthesis.
@@ 1 Introduction
This work addresses these gaps by reformulating the sampling of diffusion models and rectified flows via Stochastic Differential Equations (SDEs) and applying Group Relative Policy Optimization (GRPO) [20, 21] to stabilize the training process. GRPO is a recent breakthrough that enhances the performance of large language models (LLMs) on complex tasks such as mathematics and coding. In this paper, we pioneer the adaptation of GRPO to visual generation tasks through the DanceGRPO framework, establishing a "harmonious dance" between GRPO and visual generation tasks.
We present a comprehensive study of one algorithm, DanceGRPO across two generative paradigms (diffusion/rectified flows), three tasks (text-to-image, text-to-video, image-to-video), four foundation models (Stable Diffusion [2], HunyuanVideo [22], FLUX [23], SkyReels-I2V [24]), and five reward models (image/video aesthetics, text-image alignment, video motion quality, and binary thresholds).
@@
  • Unified and Pioneering. To the best of our knowledge, we are the first to extend GRPO to diffusion models and rectified flows, accomplishing multiple visual generation tasks within a unified RL framework. We achieve seamless integration between GRPO and visual generation tasks by carefully reformulating the SDEs, selecting appropriate optimized timesteps, and initializing noise and noise scales.
  • Generalization and Scalability. To our knowledge, DanceGRPO is the first RL-based unified framework capable of seamless adaptation across diverse generative paradigms, tasks, foundational models, and reward models. Unlike prior RL algorithms, primarily validated on text-to-image diffusion models on small-scale datasets, DanceGRPO demonstrates robust performance on large-scale datasets, showcasing both scalability and practical applicability.
@@ 2.2 DanceGRPO
$&\mathbf{s}_{t}\triangleq(\mathbf{c},t,\mathbf{z}_{t}),\quad\pi(\mathbf{a}_{t}\mid\mathbf{s}_{t})\triangleq p(\mathbf{z}_{t-1}\mid\mathbf{z}_{t},\mathbf{c}),\quad P(\mathbf{s}_{t+1}\mid\mathbf{s}_{t},\mathbf{a}_{t})\triangleq\left(\delta_{\mathbf{c}},\delta_{t-1},\delta_{\mathbf{z}_{t-1}}\right)\\ &\mathbf{a}_{t}\triangleq\mathbf{z}_{t-1},\quad R(\mathbf{s}_{t},\mathbf{a}_{t})\triangleq\begin{cases}r(\mathbf{z_{0}},\mathbf{c}),&\text{if }t=0\\ 0,&\text{otherwise}\end{cases},\quad\rho_{0}(\mathbf{s}_{0})\triangleq\left(p(\mathbf{c}),\delta_{T},\mathcal{N}(\mathbf{0},\mathbf{I})\right)\end{aligned},$ (6)
@@
3:for training iteration =1=1= 1 to $M$ do
4:  Sample batch $\mathcal{D}_{b}\sim\mathcal{D}$triangleright Batch of prompts

### v3

@@ Abstract
Recent advances in generative AI have revolutionized visual content creation, yet aligning model outputs with human preferences remains a critical challenge. While Reinforcement Learning (RL) has emerged as a promising approach for fine-tuning generative models, existing methods like DDPO and DPOK face fundamental limitations - particularly their inability to maintain stable optimization when scaling to large and diverse prompt sets, severely restricting their practical utility. This paper presents DanceGRPO, a framework that addresses these limitations through an innovative adaptation of Group Relative Policy Optimization (GRPO) for visual generation tasks. Our key insight is that GRPO’s inherent stability mechanisms uniquely position it to overcome the optimization challenges that plague prior RL-based approaches on visual generation. DanceGRPO establishes several significant advances: First, it demonstrates consistent and stable policy optimization across multiple modern generative paradigms, including both diffusion models and rectified flows. Second, it maintains robust performance when scaling to complex, real-world scenarios encompassing three key tasks and four foundation models. Third, it shows remarkable versatility in optimizing for diverse human preferences as captured by five distinct reward models assessing image/video aesthetics, text-image alignment, video motion quality, and binary feedback. Our comprehensive experiments reveal that DanceGRPO outperforms baseline methods by up to 181% across multiple established benchmarks, including HPS-v2.1, CLIP Score, VideoAlign, and GenEval. Our results establish DanceGRPO as a robust and versatile solution for scaling Reinforcement Learning from Human Feedback (RLHF) tasks in visual generation, offering new insights into harmonizing reinforcement learning and visual synthesis.
@@ 1 Introduction
This work addresses these gaps by reformulating the sampling of diffusion models and rectified flows via Stochastic Differential Equations (SDEs) and applying Group Relative Policy Optimization (GRPO) [20, 21] to stabilize the training process. In this paper, we pioneer the adaptation of GRPO to visual generation tasks through the DanceGRPO framework, establishing a "harmonious dance" between GRPO and visual generation tasks. Our key insight is that GRPO’s architectural stability properties provide a principled solution to the optimization instabilities that have limited previous RL approaches to visual generation.
We extend a systematic study of DanceGRPO, evaluating its performance across generative paradigms (diffusion models, rectified flows) and tasks (text-to-image, text-to-video, image-to-video). Our analysis employs diverse foundation models [2, 22, 23, 24] and reward metrics to assess aesthetic quality, alignment, and motion dynamics. Furthermore, through the proposed framework, we discover insights regarding the rollout initialization noise, the reward model compatibility, the Best-of-N inference scaling, the timestep selection, and the learning on binary feedback.
@@
  • Stability and Pioneering. We present the first discovery that GRPO’s inherent stability mechanisms effectively address the core optimization challenges in visual generation that have persistently hindered prior RL-based approaches. We achieve seamless integration between GRPO and visual generation tasks by carefully reformulating the SDEs, selecting appropriate optimized timesteps, initializing noise, and noise scales.
  • Generalization and Scalability. To our knowledge, DanceGRPO is the first RL-based unified application framework capable of seamless adaptation across diverse generative paradigms, tasks, foundational models, and reward models. Unlike prior RL algorithms, primarily validated on text-to-image diffusion models on small-scale datasets, DanceGRPO demonstrates robust performance on large-scale datasets, showcasing both scalability and practical applicability.
@@ 2.2 DanceGRPO
$\begin{aligned} &\mathbf{s}_{t}\triangleq(\mathbf{c},t,\mathbf{z}_{t}),\quad\pi(\mathbf{a}_{t}\mid\mathbf{s}_{t})\triangleq p(\mathbf{z}_{t-1}\mid\mathbf{z}_{t},\mathbf{c}),\quad P(\mathbf{s}_{t+1}\mid\mathbf{s}_{t},\mathbf{a}_{t})\triangleq\left(\delta_{\mathbf{c}},\delta_{t-1},\delta_{\mathbf{z}_{t-1}}\right)\\ &\mathbf{a}_{t}\triangleq\mathbf{z}_{t-1},\quad R(\mathbf{s}_{t},\mathbf{a}_{t})\triangleq\begin{cases}r(\mathbf{z_{0}},\mathbf{c}),&\text{if }t=0\\ 0,&\text{otherwise}\end{cases},\quad\rho_{0}(\mathbf{s}_{0})\triangleq\left(p(\mathbf{c}),\delta_{T},\mathcal{N}(\mathbf{0},\mathbf{I})\right)\end{aligned},$ (6)
@@
3:for training iteration =1=1 to $M$ do
4:  Sample batch $\mathcal{D}_{b}\sim\mathcal{D}$ $\triangleright$ Batch of prompts